### v1

# The Atacama Cosmology Telescope: Release of A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT)

Abstract
We present A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT) hosted on Amazon Web Services, Inc. (AWS) in the form of a public Amazon Simple Storage Service (S3) bucket. This bucket is an Amazon cloud storage database containing flux measurements for a group of asteroids at millimeter (mm) wavelengths. These measurements were collected by the Atacama Cosmology Telescope (ACT) from 2017 to 2021 in frequency bands centered near $90$, $150$, and $220$ GHz. The ASTRONAUT database contains observation times, normalized flux values, and associated error bars for 170 asteroids above a signal-to-noise ratio of 5 for a single frequency band over the stacked co-added maps. We provide an example in generating light curves with this database. We also present a Jupyter notebook to serve as a reference guide when using the S3 bucket. The container and notebook are publicly available in a GitHub repository.
Asteroids (72), CMB (322), $m$ Astronomy (1061)
1 Introduction
Studies of asteroids in the solar system have increased our understanding of early planetary formation processes as diverse asteroid spectra have indicated a dynamic solar system history (DeMeo & Carry, 2014). Studying the composition of asteroid sub-surfaces (commonly referred to as the asteroid regolith) can further increase our understanding of physical processes in the nascent solar system. Asteroid measurements at millimeter wavelengths provide a way to study regolith composition since these measurements are sourced from below the asteroid surface (Ulich & Conklin, 1976; Johnston et al., 1982; Viikinkoski et al., 2015). Different asteroid sub-surface compositions can lead to varying flux measurements as a function of wavelength. Comparing these measurements to models in the infrared (IR) have suggested a sub-millimeter (sub-mm) deficit (e.g., results from Müller & Barnes (2007); Orlowski-Scherer et al. (2024); Chichura et al. (2022)).
Changes in asteroid composition can signify the presence of temperature gradients or changes in emissivity as a function of wavelength, although more work is needed to better explain which process is responsible. Millimeter observations can therefore shed light on the formation of the asteroid belt in the solar system. However, there is a lack of millimeter flux data from asteroids compared to other wavelengths. Most asteroid studies have been performed in the optical and IR regimes, requiring large observing resources for targeted searches (Müller & Barnes, 2007; Chamberlain et al., 2007). Millimeter observations are also limited by the apparent size of asteroids, as well as their approximately black body spectral radiance profile. Additionally, asteroids are quite faint at these wavelengths compared to the optical/infrared wavelengths where most asteroid science is done. This is both due to their spectral slope and the diffraction limit imposing a $10^{3}-10^{6}$ times worse beam dilution factor. Millimeter observations are therefore relatively costly, and have until recently been limited to targeted observations of a few asteroids.
Recent studies have demonstrated the feasibility of collecting asteroid thermal emission flux from cosmology observatories (Chichura et al., 2022; Orlowski-Scherer et al., 2024). Cosmic Microwave Background (CMB) survey experiments collect data in the millimeter to sub-millimeter range, which makes telescopes such as the South Pole Telescope (SPT), the Atacama Cosmology Telescope (ACT), and the Simons Observatory (SO) well-suitedcandidates for studying asteroids in this wavelength regime (Abitbol et al., 2025). Additionally, the CCAT observatory will include a higher frequency range than ACT, while also providing overlapping sky coverage for complementary asteroid spectral information (Aravena et al., 2022). As part of their observations of the CMB, these observatories make repeated observations of large fractions of the sky. This results in incidental observations of large numbers of asteroids.
To increase access to these data, we present A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT) hosted by Amazon Web Services, Inc. (AWS). This database is in the form of a public Amazon Simple Storage Service (S3) bucket. This bucket contains normalized thermal emission flux measurements, associated error bars, weighting factors based on asteroid geometry, and observation times. To aid in the S3 bucket user implementation, we also include a Jupyter notebook, which is hosted on the ASTRONAUT GitHub repository.  [^1] https://github.com/ACTCollaboration/ASTRONAUT This notebook details methods to pull the flux data and create light curves.
In this paper, we present data products from an analysis of 170 asteroids extracted from observations made by ACT in frequency bands centered near 98, 150, and 228 GHz, denoted f090, f150, and f220, respectively (Orlowski-Scherer et al., 2024). ACT can probe into the regolith with depths of a few millimeters to a few centimeters, although these distances are influenced by different material properties. The asteroids presented in this study contain values at the $5$ sigma signal-to-noise level or higher in at least one observing band when stacking over all observations.
This paper is structured as follows. In Section 2, we provide an overview of the ACT data and briefly describe the depth-1 maps used to study asteroids. In Section 3, we summarize the asteroid data products from the S3 bucket and describe the bucket itself. We also discuss light curves, the main purpose for this container. In Section 4, we present a Jupyter Notebook which describes how to use the bucket and create an example light curve. Finally, in Section 5, we summarize this work.
2 The Atacama Cosmology Telescope
The Atacama Cosmology Telescope was a $6$ m off-axis Gregorian telescope located in the Atacama Desert in Chile (Fowler et al., 2007; Thornton et al., 2016), which was predominately used for millimeter observations of the cosmic microwave background between 2007 to 2022. ACT made CMB observations in six frequency bands: f030030, f040040, f090090, f150150, f220220, and f280280 GHz (Li et al., 2021). The database presented here contains observations of asteroids from the f090090, f150150, and f220220 GHz channels during 2017-2021. Those three bands are used because maps do not exist for $30$ and $40$ GHz, while $280$ GHz observations were done on small patches of the sky, which makes finding asteroids extremely difficult. These three frequency bands are centered around $98$, $150$, and $228$ GHz, with beam full width at half-maximum of $2.0^{\prime}$, $1.4^{\prime}$, and $1.0^{\prime}$, respectively. As ACT uses dichroic detectors, these frequency bands correspond to multiple ACT arrays: pa4, pa5, and pa6. Table 1 indicates the observing frequencies in each array.
2.1 Depth-1 Maps
ACT maps the sky by scanning at constant elevation while the sky drifts past, resulting in a relatively shallowly exposed image called a “depth-1” map in ACT Data Release 6, which forms the basis of this analysis (Naess et al., 2025). One can think of depth-1 maps as single-exposure images with an exposure time of around 5 minutes. However, they are not technically single-exposure, as the instrument is continuously read out, and the telescope makes multiple azimuth swipes across a given spot in the sky while it drifts past. The depth-1 maps are made using a modified version of the maximum-likelihood procedure used for the standard ACT maps with fewer conjugate gradient steps (Naess et al., 2020). This saves computational resources at the cost of biasing the maps on large angular scales irrelevant to the point source nature of asteroids. Depth-1 maps are made for each array and associated frequencies of observation.
2.2 Data Processing
We measure the asteroid fluxes from the depth-1 maps following Orlowski-Scherer et al. (2024). The depth-1 maps are matched filtered for point sources, which cuts out atmospheric noise scales as well as scales smaller than the beam size. This matched-filter maximizes the signal-to-noise (S/N) for point sources on the maps, which includes asteroids. Orlowski-Scherer et al. (2024) then create small stamps of depth-1 maps using the JPL Horizons API. [^2] https://ssd.jpl.nasa.gov/horizons/ This API was implemented to determine the location of asteroids on the depth-1 maps using the ACT observation times. Using the Horizons ephemerides, it was possible to interpolate the orbit of asteroids provided observing times. Once these asteroids are centered, measuring the flux corresponds to evaluating the matched-filter flux map value at the location of the asteroid on a single depth-1 map. The flux error is calculated from a flux uncertainty map. These flux values can be used to create light curves, as described in Section 3.2. This database therefore contains the flux values returned from the center of depth-1 stamps after a normalization procedure, described below. This also includes the flux error, as well as the weighting values used to derive the normalized fluxes. The time corresponds to the observation time of the asteroid, up to the array crossing time (approximately $4$ minutes), which is returned in Unix time units. Unix time is a unit of measurement corresponding to the number of non-leap seconds that have passed since 00:00:00 UTC on January 1, 1970 (Stevens & Rago, 2013). We refer readers to Orlowski-Scherer et al. (2024) for more details.
Figure 1: Example plot of the depth-1 maps discussed in Section 2.1 which are used to generate light curves discussed in Section 3.2. Light curve data shown here are for (4) Vesta on array pa5 at $150$ GHz. The color bar corresponds to flux intensity. Note that this data release does not contain data maps such as those shown here.
3 Data Products
For each ACT array and frequency band, we provide normalized flux data, flux-errors, weighting factors, and time of observation in the public S3 database. The flux is normalized by the expected flux under the Rayleigh-Jeans limit of the Standard Thermal Model (Lebofsky et al., 1986). We normalize this flux due to the varying asteroid and Earth/Sun distances between depth-1 maps, as well as the changing observing angle. We can better compare or combine measurements at different geometries through this normalization process. The Rayleigh-Jeans limit can be applied due to the relatively simple spectral energy distribution of asteroids in the far-infrared range (Mok et al., 2003). We catalog the flux as follows:
$F_{0}$ $=F_{i}(\frac{d_{earth,i}}{1AU})^{2}(\frac{d_{sun,i}}{1AU})^{1/2}10^{0.004\alpha_{i}}$ (1)
$=F_{i}W_{i}$ (2)
where the weighting scheme $W_{i}$ is based on the orbital position of the asteroid, which includes the Earth $d_{earth,i}$ and Sun $d_{sun,i}$ centered distances, as well as the Earth-asteroid-Sun angle $\alpha_{i}$ in degrees. $F_{0}$ represents the normalized flux for $d_{earth}=d_{sun}=1$ AU. We apply this normalization procedure for each observation stamp $i$. The associated error bars are also normalized in this fashion. The database therefore contains $F_{0}$, $W_{i}$, flux error, and observation time. We refer readers to Table 2 in the Appendix for an overview of the format of the data.
In addition to using the S3 bucket, these data can also be accessed on the Legacy Archive for Microwave Background Data Analysis (LAMBDA) webpage.  [^3] https://lambda.gsfc.nasa.gov/ The ASTRONAUT data on LAMBDA will be in the form of $\mathtt{tar}$ files. The $\mathtt{tar}$ files provide an alternative to querying this public bucket at the expense of less user-specific file outputs.
Table 1: Table of ACT array/frequency combinations and array sensitivities (Naess in prep.) supported by this database.
ACT Array Frequency (GHz) Sensitivity ($\mu K\sqrt{s}$)
pa4 $150,220$ $19.1,42.6$
pa5 $90,150$ $12.6,15.1$
pa6 $90,150$ $11.4,13.6$
3.1 Amazon Simple Storage Service
A databaSe of millimeTeR ObservatioNs of Asteroids Using acT (ASTRONAUT) is maintained on a public S3 bucket hosted by Amazon’s cloud storage service and is located on Amazon’s registry of open data on AWS. [^4] https://registry.opendata.aws/hst/ The S3 bucket is implemented in a python script. Data can be accessed via python scripts and/or Jupyter Notebooks.
The bucket returns flux, flux error, weights, and observation times in the form of a $\mathtt{FITS}$ file named $\mathtt{name\_lc\_arr\_freq.fits}$ for each specific query, with $\mathtt{name}$, $\mathtt{arr}$, and $\mathtt{freq}$ being user-defined.
Since the ACT arrays are dichroic, not all frequencies are supported by each array. Therefore, certain queries to the S3 bucket will result in client errors if care is not taken (for example, requesting data on array pa6 for frequency $220$ GHz). A result of an invalid request will return $\mathtt{ClientError}$ to the user.
3.2 Light Curves
By querying this S3 bucket for a specific asteroid, ACT array, and frequency, users can create their own light curve plots. We provide an example plot for (705) Erminia shown in Figure 2. This plot can be recreated using the code described in the Notebook tutorial.  [^5] https://github.com/ACTCollaboration/ASTRONAUT/blob/main/Notebooks/Tutorial/Tutorial.ipynb
Figure 2: Example light curve at 90 GHz for (705) Erminia using the ASTRONAUT database. Partial data can be seen in Table 2 of the Appendix.
We discourage users from comparing asteroids based on their time-averaged signal. Specifically, averaging across different asteroid depth-1 maps will not yield interpretable results due to differences in regolith composition. To average the data, we recommend placing the asteroids into specific classes, which is discussed in Orlowski-Scherer et al. (2024).
3.3 Available Asteroids
We also create a script for user-specific requests. For example, users may want a specific asteroid in the database, but may not know if it exists. Using our script with appropriate functions, users may query the $\mathtt{asteroid\_list.txt}$ file for specific requests. We also include functions to search for specific array and frequency combinations. We provide instructions on how to use the script in a markdown file. [^6] https://github.com/ACTCollaboration/ASTRONAUT/blob/main/lookup.md
4 Jupyter Notebook
In addition to launching this database, we also create a Jupyter Notebook to aid users. This Notebook guides users through requesting data from the public bucket (described in Section 3), reading in the data file(s), and generating an example light curve.
This notebook can be accessed on the ASTRONAUT GitHub site.  [^7] https://github.com/ACTCollaboration/ASTRONAUT/blob/main/Notebooks/Tutorial/Tutorial.ipynb The main purpose of the GitHub is to provide more specific information for queries and host the Jupyter Notebook tutorial. We also provide a file for users to download necessary dependencies.
4.1 Notebook Tutorial
As an example, the notebook describes how to generate asteroid light curves. It starts with users installing relevant python packages, which can be found in the $\mathtt{requirements.txt}$ file in the ASTRONAUT GitHub. We then describe how users make data requests for specific asteroid(s) on different arrays at different frequencies. We currently support data requests to multiple arrays and frequencies for a specific asteroid. This allows users to generate a light curve across more than one frequency. We then display an example light curve in the Notebook, shown in Figure 2.
5 Conclusion
In this paper, we describe a database containing time, normalized flux, flux uncertainty, and weighting factors for 170 asteroid observations in the millimeter wavelengths. We present this database in the form of a public Amazon Simple Storage Services bucket. These data were collected by the Atacama Cosmology Telescope from 2017-2021. We describe general features of the container, such as data products and basic implementation. We also present a tutorial for users, which can be accessed via Jupyter Notebooks on a GitHub repository. These data are freely available and we encourage the use of these data.
Future applications of these data include creating light curves across observatories, such as ACT, SPT, and the Atacama Large Millimeter/submillimeter Array, as well as calculating phase curves to better understand the variation of millimeter flux as a function of asteroid rotation. These data could also be used to calibrate millimeter instruments such as those mentioned above.
6 Acknowledgments
This work was supported by the U.S. National Science Foundation through awards AST-1440226, AST0965625 and AST-0408698 for the ACT project, as well as awards PHY-1214379 and PHY-0855887. The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. ACT operates in the Parque Astronómico Atacama in northern Chile under the auspices of the Comisión Nacional de Investigación Científica y Tecnológica de Chile (CONICYT), now La Agencia Nacional de Investigación y Desarrollo (ANID). Colleagues at AstroNorte and RadioSky provide logistical support and keep operations in Chile running smoothly.Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto.
This work was supported by a grant from the Simons Foundation (CCA 918271, PBL). RCV acknowledges funding and support from the Nexus Scholars Program. Nick Battaglia acknowledges the support from NASA grants 21-ADAP21-0114 and 21-ATP21-0129.
We thank Bez Thomas for their technical support.